How many milligrams of sodium chloride may be used in the preparation of 15 mL of an eye drop containing 1% tropicamide and 0.5% chlorobutanol to render the solution isotonic with tears?
Answer #1

Solution: Given, 15 mL of an eye drop preparation,

1% tropicamide means 100ml contains 1gram.

1 × 0.21 = 0.21 g of NaCl represented by Tropicamide.

So, 100 × 0.009 = 0.9 g of NaCl in 100 mL of isotonic solution

Therefore, 0.9 - 0.21 = 0.69 g of NaCl required to make isotonic solution

0.69 g ×15 ml required to be prepared = 10.35 gram

So, 10.35 × 10 = 103.5mg sodium chloride.